OpenStudy (anonymous):

How to do epsilon-delta proof of lim_(x->0) (x^2-2x+3)=3?

OpenStudy (anonymous):

you want to show that given any ϵ>0 you can find δ so that if |x−3|<δ you get |x2−2x+3−3|<ϵ as usual we work backwards

OpenStudy (anonymous):

|x−0|<δ?

OpenStudy (anonymous):

\[|x^2-2x+3-3|=|x^2-2x|=|x||x-2|\] and we get to control |x|

OpenStudy (anonymous):

yes i was being stupid i know, but i just wanted it to look like \[|x-a|<\delta\] in this case of course \(a=0\)

OpenStudy (anonymous):

now since we control |x| by making it as small as we like, we only have to get a handle on |x-2|

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

so the usual trick is since x is near zero we can say assume that |x|<1 so |x-2|<3

OpenStudy (anonymous):

|x-2| < C

OpenStudy (anonymous):

|x-2| < 1 and -1 < x < 1 -3 < x-2 < -1

OpenStudy (anonymous):

this line was wrong i meant \[|x||x-2|< \frac{\epsilon}{3}\times 3=\epsilon\]

OpenStudy (anonymous):

satellite are u sure that |x - 2|<3 ? Cuz, if |X|<1 so |x-2|< 1-2 -> |x-2|< -1

OpenStudy (anonymous):

let me write it again correctly assume |x|<1 so -1<x<1 and therfore -3<x-2<-1 ensuring |x-2|<3

OpenStudy (anonymous):

we want an upper bound for the term |x-2| and i got 3 you can of course find a different one but now that i have bounded |x-2|<3 i can pick \(\delta=\frac{\epsilon}{3}\)

OpenStudy (anonymous):

therefore is \(|x|<\frac{\epsilon}{3}\) and \(\frac{\epsilon}{3}<1\) we get \[|x||x-2|<\frac{\epsilon}{3}\times 3=\epsilon\]

OpenStudy (anonymous):

satellite, is it honest?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i'm talking about mathematical honesty in this part: assume |x|<1 so -1<x<1 and therfore -3<x-2<-1 ensuring |x-2|<3 I'm confused